Solve each of the following systems by substitution.
5x-2y=-5
y-5x=3

Answer by HyperBrain(694)

y-5x=3
___y=5x+3
5x-2y=-5
But,
___y=5x+3
thus,
5x-2(5x+3)=-5
__5x-10x-6=-5
_______-5x=1
_________x=-1/5
Solving for y,
y=5(-1/5)+3
_=-1+3
_=2
Therefore, x==-1/5 and y=2
Power up,
HyperBrain